A person's level of blood glucose and diabetes are closely related. Let x be a random variable measured in milligrams of glucose per deciliter (1/10 of a liter) of blood. Suppose that after a 12-hour fast, the random variable x will have a distribution that is approximately normal with mean μ = 73 and standard deviation of σ = 20. What is the probability that, for an adult after a 12-hour fast, x is more than 85?

Select one:

a. 0.726

b. 0.274

c. 0.363

d. 0.226

e. 0.113
Answer #1

μ = 73
σ = 20
X = 85

P(X > 85) = P(Z > (x - μ)/σ) = P (Z > (85 - 73)/20) = P(Z > 0.6) = 0.2743 (from z table)

correct option is B

please let me know if this helped